OpenStudy (anonymous):

Evaluate the following indefinite integral by using substitution: \[\int\limits_{}^{}(\sin(2x+10)/\cos^2(2x+10))dx\]

OpenStudy (anonymous):

u=cos(2x+10) and see what happens

OpenStudy (amistre64):

\[-\frac{1}{2cos(2x+10)}\] maybe

OpenStudy (amistre64):

-2+1 = -1 so im sure that parts right lol

OpenStudy (anonymous):

ghass: why can you set u to equal that? i thought u has to be something that is "inside a function"

OpenStudy (anonymous):

let \[2x+10=t\] implies dx=dt/2 \[ \int\limits sint/2\cos ^{2}t dt\] solve it basic integrations

OpenStudy (amistre64):

\[\frac{1}{[...]}\] is the inside of a function

OpenStudy (anonymous):

ok so i actualy got what luckey got

OpenStudy (anonymous):

should the next step be a substitution as well?

OpenStudy (amistre64):

all subbing does is "clean things up" so they are easier to see

OpenStudy (anonymous):

u is a function of x u=cos(2x+10) du=-2sin(2x+10)dx... then your integral becomes (-1/2)*int(du/u^2) which is easy 2 compute

OpenStudy (anonymous):

ah ok i see it ty guys

OpenStudy (amistre64):

easy versus familiar ...

OpenStudy (turingtest):

I got positive amistre's answer\[-{1\over2}\int\limits {du \over u^2}={1\over2u}+C={1\over2\cos(2x+10)}+C\] Ghass' techinique is correct

OpenStudy (anonymous):

then read it as \[1/2 \int\limits \tan(t) \sec(t)dt\] which is equal to \[1/2sect\] substitute the value of t u get 1/2sec(2x+10)

OpenStudy (amistre64):

accursed signs lol

OpenStudy (anonymous):

i got the same answer as test got........

OpenStudy (turingtest):

yes and used a very different method :)